OpenStudy (anonymous):

if g/h = 6/5 which equation is true

OpenStudy (anonymous):

What are the equations?

OpenStudy (anonymous):

a)5h=6g b) h/g=5/6 c) h/6=g/5 gh=6x5

OpenStudy (anonymous):

B is right.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Click 'Best Response' next to my comment to give me a medal!

OpenStudy (anonymous):

6/a = 18/27 solve the proportion: a)54 b) 81 c)9 d)18

OpenStudy (anonymous):

C